Over the past few decades, Boeing’s chief competitor was Airbus. Recently, smaller firms have entered the market. One of these firms is Embraer. Embraer recently had one of the best-selling business jets, the Phenom 300. Though, there are competitors, Boeing has led the market in terms of pricing.

1. What type of market structure is this and what model might we use to analyze it? These companies have the following total cost functions with A = Airbus, E = Embraer and B = Boeing: TCA= $1,500,000 + $30,000,000QA + $500,000QA2 TCE= $484,000,000 + $10,000,000QE + $250,000QE2 TCB= $3,000,000,000 + $2,000,000QB + $55,000QB2 The industry demand curve for this type of jet aircraft is: Q = 950 - 0.000015P

2. What are the supply functions for Airbus and Embraer? (Hint: these firms operate as price takers)

3. What is Boeing’s demand function? (Hint: Boeing’s demand function is the industry demand curve minus the following firms’ supply or QB= Q – QA– QE. Use the answers you found in question 2)

4. What is Boeing’s profit-maximizing price and output.

5. What are Airbus’ and Embraer’s profit-maximizing output levels?

6. Is the industry in short-run equilibrium, i.e. does firm supply equal total demand?

7. Which firms are earning an economic profit? Which ones are earning a normal rate of return? How do you know?

8. Is the industry in a long-run equilibrium? Why or why not?

I have answered 1-4 but need help with 5-8

Can you have it done by February 20th?
Answer #1

as there are small number of sol (0 the market companies structure controling is oligopolistic the market The profit marimisa= 2m +0.11QB Equating this price am to 11m QA = 1000 - QB 0.000018 36 +198 Q1 = 100000 OR ² 323 323 1000 - 0000018p P = 37.58